OpenStudy (anonymous):

What is the discriminant of x2 - 7x + 2 = 0? (3 points) A) -1 B) 41 C) 0 D) -57

OpenStudy (anonymous):

the discriminant is given by b^2 - 4ac a = 1 b = -7 What is c?

jhonyy9 (jhonyy9):

D=49-8=41 B. is right